OpenStudy (anonymous):

I'm trying to solve the quadratic equation f(x)= -2(x +2)^2 but I don't know how to use the form ax^2+bx+c without the +c, can someone explain? please

OpenStudy (dumbcow):

right now its in vertex form you have to expand to get it into standard form

OpenStudy (anonymous):

I have to find the vertex

OpenStudy (dumbcow):

(x+2)^2 = (x+2)(x+2) distribute

OpenStudy (dumbcow):

oh nevermind, then its easy

myininaya (myininaya):

you have it in vertex form \[f(x)=a(x-h)^2+k, V(h,k)\]

myininaya (myininaya):

\[f(x)=-2(x+2)^2+0\]

myininaya (myininaya):

identify what h and k are and you are done

OpenStudy (anonymous):

ok, just make it 0, it's that easy, thank you